OpenStudy (anonymous):

in square roots, when you add them, do you reduce the final answer? and when you add what if the number inside isnt the same ans the other one? ex. sqrt of 3 + sqrt of 6

OpenStudy (anonymous):

you cannot add the square roots of two different numbers, so your answer cannot be reduced. \[\sqrt{3} + \sqrt{6}\] is unsolvable if you had numbers that are square roots or multiples of square roots, you can reduce the numbers to add them. you could, for example add \[3\sqrt{3} + \sqrt{3} = 4 \sqrt{3}\]

OpenStudy (anonymous):

sooo... if i had -3\[\sqrt{2} - 2\sqrt{8}\]

OpenStudy (anonymous):

i would reduce the -8 so it would be 4x2

OpenStudy (anonymous):

\[\sqrt{8} = \sqrt{2^2\cdot 2} = 2 \sqrt{2}\]\[\implies 2\sqrt{8} = 2(2\sqrt{2}) = 4\sqrt{2}\]

OpenStudy (anonymous):

then he answer would be -\[\sqrt{2}\]

OpenStudy (anonymous):

you can reduce \[2\sqrt{8}\] to \[4\sqrt{2}\]

OpenStudy (anonymous):

So you have one \(\sqrt{2}\) and you subtract 4 \(\sqrt{2}\)s. Leaves you with -3 \(\sqrt{2}\)s = \(-3\sqrt{2}\)

OpenStudy (anonymous):

so it would be -sqrt2

OpenStudy (anonymous):

think of it as x - 4x, its -3x or in this case \[-3\sqrt{2}\]